OpenStudy (anonymous):

write an equation for each tangent to the curve y = 1 / (x - 1) that has slope -1

OpenStudy (anonymous):

take the derivative, set it equal to one, solve the quadratic

OpenStudy (anonymous):

1 or -1?

OpenStudy (anonymous):

\[-\frac{1}{(x-1)^2}=-1\] \[(x-1)^2=1\] etc

OpenStudy (anonymous):

oh sorry i meant \(-1\) as you wrote

OpenStudy (anonymous):

\[(x-1)^2=1\] \[x-1=1\implies x=2\] or \[x-1=-1\implies x=0\]

OpenStudy (anonymous):

ok. and now i need to write an equation for each